### v1

# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

Abstract
Chain-of-Thought (CoT) prompting has significantly advanced the reasoning capabilities of Large Language Models (LLMs), yet the reliability of these reasoning chains remains a critical challenge. A widely held “cascading failure” hypothesis suggests that errors are most detrimental when they occur early in the reasoning process. This paper challenges that assumption through systematic error-injection experiments, revealing a counter-intuitive phenomenon we term “Late-Stage Fragility”: errors introduced in the later stages of a CoT chain are significantly more likely to corrupt the final answer than identical errors made at the beginning. To address this specific vulnerability, we introduce the Adaptive Self-Correction Chain-of-Thought (ASCoT) method. ASCoT employs a modular pipeline in which an Adaptive Verification Manager (AVM) operates first, followed by the Multi-Perspective Self-Correction Engine (MSCE). The AVM leverages a Positional Impact Score function $I(k)$ that assigns different weights based on the position within the reasoning chains, addressing the Late-Stage Fragility issue by identifying and prioritizing high-risk, late-stage steps. Once these critical steps are identified, the MSCE applies robust, dual-path correction specifically to the failure parts. Extensive experiments on benchmarks such as GSM8K and MATH demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines, including standard CoT. Our work underscores the importance of diagnosing specific failure modes in LLM reasoning and advocates for a shift from uniform verification strategies to adaptive, vulnerability-aware correction mechanisms.
Introduction
The advent of Chain-of-Thought (CoT) has endowed Large Language Models (LLMs) with remarkable multi-step reasoning capabilities (Nye et al. 2021; Kojima et al. 2022), enabling them to deconstruct complex problems into a sequence of manageable sub-problems (Yao et al. 2023; Shinn et al. 2023). Recent studies, such as the o1 from OpenAI  (Jaech et al. 2024) and DeepSeekR1  (Guo et al. 2025) have shown that extended CoT reasoning improves LLM performance. However, the success of CoT and its more verbose variants has introduced challenges: prohibitive computational costs from long reasoning chains (Chen et al. 2025) and, more critically, the inherent brittleness of the reasoning process itself (Zhang et al. 2024; Ke et al. 2025). A single flaw is often sufficient to invalidate an entire process of thought, raising significant concerns about the trustworthiness of these models (Chowdhury and Caragea 2025).
Figure 1: The top part (a) illustrates the standard CoT reasoning process, which often generates redundant outputs and lacks an effective self-correction mechanism when errors occur. The bottom part (b) depicts the ASCoT method, which incorporates a robust mechanism for self-correction, addressing potential reasoning flaws at each stage.
A widely held belief within the research community, often termed the “cascading failure” hypothesis, posits that the primary threat to CoT integrity lies in early-stage errors (Tyen et al. 2023). The logic is intuitive: an initial mistake in problem comprehension or a miscalculation will inevitably propagate through the logical dependency graph of the reasoning chain, contaminating all subsequent steps. Consequently, much of the research on improving CoT robustness has implicitly focused on ensuring the correctness of the initial steps through more robust prompting or post-hoc verification (Liu et al. 2024c; Yi, Wang, and Li 2025).
In this work, we directly investigate this foundational assumption, seeking to answer a critical question: Are all errors in a CoT created equal in their impact? To find out this, we conduct a systematic error-injection experiment, which introduces controlled symbolic and numeric errors at varying positions within correct reasoning chains. We then instruct the LLM to resume the reasoning process from the point at which the error occurred. Our findings are striking and counter-intuitive. We discover that errors introduced in later stages are more likely to lead to an incorrect final answer. We term this phenomenon “Late-Stage Fragility”. An early error often appears to trigger the model’s latent self-correction mechanisms, allowing it to recover and still reach the correct solution. However, as the model progresses steadily, its reasoning process gains a form of “semantic commitment”, reducing relevant information entropy and flexibility. At the later stages, the model is unable to evaluate the validity of the final calculations.
This empirical discovery suggests that a significant portion of efforts to secure CoT reasoning may be misdirected. As illustrated in Figure  1(a), a standard CoT process can be both inefficient due to redundancy and unreliable due to its inability to correct these late-stage errors. Building on a rethinking of how we evaluate and how we correct LLM reasoning, we introduce the ASCoT method. In contrast to standard methods, ASCoT is designed to address both challenges simultaneously. As shown in Figure  1(b), it first employs a semantic pruning stage to enhance efficiency, followed by a targeted, risk-aware correction mechanism that focuses computational resources precisely on the most vulnerable steps to ensure a correct final answer.
We evaluate ASCoT across various models, including the LLaMA-3.1-8B-Instruct and the Qwen2.5-Instruct series (3B, 7B, and 14B), using two widely recognized reasoning benchmarks: GSM8K and MATH.
To sum up, our contributions are summarized as follows:
  • As far as we know, this is the first to identify and quantify the “Late-Stage Fragility” phenomenon in CoT reasoning, demonstrating that errors in later steps are significantly more detrimental than those in earlier steps.
  • We introduce ASCoT, a novel method specifically designed to counteract this fragility by precisely identifying and robustly correcting the high-impact, late-stage errors that are most critical to reasoning accuracy.
  • Our extensive experiments validate the effectiveness of ASCoT. We demonstrate that ASCoT achieves outstanding accuracy on the GSM8k and MATH benchmarks, validating our approach of diagnosing the specific failure mode before engineering a targeted solution.
Preliminary
This section establishes the foundational concepts that motivate our work, offering a review of some effective methods for identifying token importance within CoT.
Efficient CoT in LLM Reasoning
A critical question for CoT efficiency is whether every token in different position contributes equally to the final answer (Xia et al. 2025). Several early approaches (Hou et al. 2022; Zhang et al. 2023; Chen et al. 2024a) measure importance based on the self-information or surprisal from a unidirectional language model:
$I_{1}(x_{i})=-\log P(x_{i}|x_{<i};\theta_{\mathcal{M}_{L}}),\quad$ (1)
where $x_{i}$ is a given token, $\theta_{\mathcal{M}_{L}}$ denotes the LLM used to compute the confidence and $P(\cdot\mid\cdot;\theta_{\mathcal{M}_{L}})$ is the conditional probability assigned by the causal LLMs. This metric suffers from known limitations, such as positional dependency, where tokens at the end of sentences often receive lower importance scores due to the nature of perplexity in LLMs.
To address this, more advanced methods like LLMLingua-2 (Pan et al. 2024) utilize a bidirectional language model (Jawahar, Sagot, and Seddah 2019; Devlin et al. 2019) (e.g., BERT-style) trained specifically for token classification. This approach measures importance by the predicted probability of a token being essential (Achiam et al. 2023), conditioned on the entire sequence:
$I_{2}(x_{i})=P(x_{i}|x_{<n};\theta_{\mathcal{M}_{L}}),\quad$ (2)
where $\theta_{\mathrm{\mathcal{M}_{L}}}$ represents the parameters of the trained bidirectional model. Such methods confirm that tokens within a CoT sequence are highly variable in importance. For example, tokens comprising mathematical equations or key entities (Ma et al. 2025) are identified as critical, whereas semantic connectors such as “so” and “since” contribute much less to the final outcome of the reasoning  (Hou et al. 2025). This highlights a significant potential for enhancing CoT efficiency by pruning these redundant tokens.
Error Propagation in Reasoning Process
Beyond efficiency, the primary challenge for CoT is reliability. A single error can invalidate an entire reasoning process (Guo et al. 2017; Tyen et al. 2023). The conventional understanding of this problem is rooted in the “cascading failure” hypothesis, which posits that errors are most damaging when they occur early in the reasoning chain (Jiang et al. 2025; Neumann and Singh 2025). According to this view, an early mistake propagates through the logical dependency graph, contaminating all subsequent steps and ensuring an incorrect outcome. In this study, our preliminary investigations reveal a counter-intuitive dynamic that challenges this assumption. Through systematic error-injection experiments, we observe that errors introduced in the later stages of a reasoning chain are significantly more detrimental to the final outcome than identical errors introduced at the beginning. We term this phenomenon “Late-Stage Fragility”.
Method
Figure 2: Illustration of ASCoT Pipeline. ASCoT generates the $CoT_{initial}$ from the target LLM, which is then compressed by the IRM module to a predefined ratio $\gamma$. The compressed CoT is verified by the AVM module, and if the confidence score exceeds a threshold $\tau$, the problematic steps are sent to the MSCE module for error correction. The corrected $CoT_{final}$ is then used for fine-tuning. Finally, ASCoT enables reasoning on new problems with the specified compression ratio $\gamma$.
In this section, we detail the architecture of our method. ASCoT generates the $CoT_{initial}$ based on the target LLM. The $CoT_{initial}$ is then processed through the IRM module, reducing the output to a fixed ratio $\gamma$. The remaining CoT is passed through the AVM module, which performs a rigorous verification process. If the model’s risk score for any CoT step exceeds a predefined threshold $\tau$, indicating a potential error, the corresponding CoT step is forwarded to the MSCE module. The MSCE module applies internal and external validation techniques to correct errors, resulting in the corrected $CoT_{final}$. Ultimately, the refined, non-redundant $CoT_{final}$ is used for fine-tuning. We illustrate the training and inference process in Figure  2, respectively.
Intelligent Routing Mechanism
The IRM serves as the initial processing stage, responsible for preliminary CoT pruning and compression. The core idea is inspired by methods TokenSkip (Xia et al. 2025), which recognize that not all tokens in a CoT are equally important for the final reasoning outcome. The mechanism operates by first generating a complete CoT path for a given problem. Then it employs a trained token importance model to assign a semantic importance score to each token in the chain. Based on a predefined compression ratio $\gamma$, tokens with scores below the corresponding threshold are pruned, resulting in a compressed CoT, namely the final answer.
Figure 3: The details of verification and self-correction. AVM and MSCE working together to correct a faulty step $t_{k}$. AVM computes $R(t_{k})$. MSCE then applies intrinsic and extrinsic correction to produce the final corrected $CoT_{final}$.
Adaptive Verification Manager
AVM dynamically identifies which steps are most likely to be incorrect and then computes a Risk Score ${R}(t_{k})$ for each step $t_{k}$ by integrating two key assessments: the step’s confidence assessment $Q(t_{k})$ and its positional impact $I(k)$.
Confidence Assessment Score
To provide a holistic measure of the single reasoning step’s quality, we first define a comprehensive quality score $Q(t_{k})$, which is a composite metric that evaluates a step $t_{k}$ across four distinct dimensions: Logical Validity $V(t_{k})$, Factual Support $G(t_{k})$, Semantic Clarity $C(t_{k})$ and Process Utility $U(t_{k})$.
Logical validity $V(t_{k})$ evaluates the internal logical consistency of step $t_{k}$. The score is a binary value that indicates if the step $t_{k}$ is derivable from its context $(P,S_{<k})$, determined by a verification query to the LLM $\mathcal{M}_{L}$:
$V(t_{k})=\mathbb{I}(o_{k}^{v}\in y),$ (3)
where $o_{k}^{v}$ is the model’s output of step $t_{k}$, $y$ is the set of valid responses and $\mathbb{I}(\cdot)$ is the indicator function.
Factual support $G(t_{k})$ measures the correctness of all $n$ verifiable steps:
$G(t_{k})=\prod_{i=1}^{n}T(a_{i}),$ (4)
where $T(a_{i})\rightarrow\{0,1\}$ is the verification function that evaluates whether the $i$-th action $a_{i}$ correctly invokes the appropriate external tool.
Semantic clarity $C(t_{k})$ quantifies the clarity of step $t_{k}$ based on the average conditional probability of its tokens $\{y_{1},...,y_{m}\}$, conditioned on the prior context:
$C(t_{k})=\frac{1}{m}\sum_{i=1}^{m}\log P(y_{i}|c_{<i}),$ (5)
where $y_{i}$ is the $i$-th token of step $t_{k}$, $m$ is the total number of tokens, $c_{<i}$ is the preceding context of $y_{i}$ and $P(\cdot\mid\cdot)$ is the conditional probability produced by the LLM.
Process utility $U(t_{k})$ measures the contribution of step $t_{k}$ to the final answer $\mathcal{A}^{*}$. It is defined as the information gain, measured by the increase in the probability of the model generating $\mathcal{A}^{*}$ when $t_{k}$ is incorporated into the context:
$U(t_{k})=\log P(\mathcal{A}^{*}|S_{\leq k})-\log P(\mathcal{A}^{*}|S_{<k}),$ (6)
where $\mathcal{A}^{*}$ is the ground‐truth answer, $S_{\leq k}$ is the full context including step $t_{k}$ and $S_{<k}$is the context excluding step $t_{k}$.
Based on these dimensions, we design the single-step comprehensive quality scoring function $Q(t_{k})$ to integrate the multi-dimensional evaluation into a quantitative score. The function’s core design principle is a veto mechanism, implemented through a multiplicative structure. Logical validity and factual support are foundational. If a step fails on either of these criteria, it should be considered invalid regardless of its other attributes. Other assessment criteria play an auxiliary role. Therefore, the function is defined as:
$Q(t_{k})=V(t_{k})\,G(t_{k})\,\Bigl{(}w_{c}\,\frac{C(t_{k})}{C_{\max}}+w_{u}\,U(t_{k})\Bigr{)},$ (7)
where $w_{c}$ and $w_{u}$ are hyperparameters that represent the weight of semantic clarity and the weight for low process utility. Figure  3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
Positional Impact Score
To model “Late-Stage Fragility”, we introduce the Position Impact Function $I(k)$, which quantifies the influence of position on the system’s stability as it situates the different-stage regime:
$I(k)=w_{a}\cdot e^{\alpha(k-1)},$ (8)
where $k$ is the step’s position, $w_{a}$ is a base impact coefficient, and $\alpha$ is a propagation rate. These parameters are empirically fitted through controlled error-injection experiments. The detailed discussion of error propagation via controlled injection can be found in the experiment section. This procedure involves systematically injecting single errors at different steps $k$ in a set of baseline correct reasoning chains and measuring the probability of final answer failure. The function $I(k)$ is then fitted to these empirical data points. Finally, AVM combines these factors into a single Risk Score:
${R}(t_{k})=I(k)\,\times\bigl{(}1-Q(t_{k})\bigr{)}.$ (9)
This score is then compared against a tunable system threshold $\tau$. If $R(t_{k})>\tau$, the step is flagged as high-risk and passed to the MSCE for correction.
Multi-Perspective Self-Correction Engine
Once a step $t_{k}$ is flagged by the AVM, the MSCE is activated to perform a robust and efficient correction. To counteract a model struggles to correct its own mistakes when using a single line of reasoning, the MSCE employs a dual-path correction strategy: Intrinsic Correction (Self-Reflection), the model is presented with the context up to step $s_{(k-1)}$ and its own flawed step $t_{k}$, and is explicitly prompted to review and amend it. Extrinsic Correction (Independent Generation), the model is given the context up to $s_{(k-1)}$ but without the flawed step $t_{k}$. It is then prompted to generate a new subsequent step from this “clean slate”.
Experiment
Experimental Setup
Models and Datasets
We primarily evaluate our method using the LLaMA-3.1-8B-Instruct (Ge et al. 2023) and Qwen2.5-Instruct series (3B, 7B, 14B) (Team 2024)models. Our evaluation leverages two notably standard math reasoning benchmarks: GSM8K (Cobbe et al. 2021), which consists of 8.5K grade school math problems requiring 2-8 reasoning steps, and MATH (Hendrycks et al. 2021), a more challenging dataset of 12.5K competition-level math problems. Given the significant computational cost associated with the MATH dataset, we follow recent research practices by evaluating our method on a carefully selected subset, MATH-500 (Lightman et al. 2023), which includes a representative sample of 500 problems.
Implementation Details
The ASCoT is implemented in PyTorch 1.13.0 and all experiments are conducted on a server with four NVIDIA GeForce RTX 3090 GPUs (24GB). For the IRM module, we use LLMLingua-2 (Pan et al. 2024) as the token importance metric to generate compressed CoT data for fine-tuning. To optimize the fine-tuning process, we adopt LoRA (Hu et al. 2022) for all fine-tuning procedures, setting the rank $r=8$ and scaling parameter $\alpha=16$ to ensure efficient training. Additional implementation details and configurations can be found in the Appendix.
Methods Ratio GSM8K MATH-500
Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio Accuracy $\uparrow$ Tokens $\downarrow$ ActRatio
Original 1.0 86.2(0.0↓) 213.17 1.00 48.6(0.0↓) 502.60 1.00
Prompt 0.9 84.1(2.1↓) 226.37 1.06 48.6(0.0↓) 468.04 0.93
0.7 84.9(1.3↓) 209.39 0.98 48.4(0.4↓) 472.13 0.94
0.5 83.7(2.5↓) 188.82 0.89 47.8(0.4↓) 471.11 0.94
Truncation 0.9 70.2(26.0↓) 202.06 0.95 47.8(0.8↓) 440.33 0.88
0.7 25.9(60.3↓) 149.97 0.70 45.0(3.6↓) 386.89 0.77
0.5 7.0(79.2↓) 103.69 0.49 27.4(21.2↓) 283.70 0.56
ASCoT 1.0 86.9(0.7↑) 214.24 1.00 48.8(0.2↑) 478.80 0.95
0.9 86.1(0.1↓) 197.65 0.92 47.6(1.0↓) 427.15 0.84
0.8 85.1(1.1↓) 169.96 0.79 47.2(1.4↓) 376.72 0.74
0.7 83.7(2.5↓) 151.32 0.71 46.8(1.8↓) 334.26 0.67
0.6 81.6(4.6↓) 128.38 0.60 41.8(6.8↓) 309.69 0.61
0.5 79.5(6.7↓) 118.57 0.55 40.1(8.5↓) 288.94 0.57
Table 1: Experimental results of ASCoT on LLaMA-3.1-8B-Instruct. We report Accuracy, average CoT token count (Tokens) and actual compression ratio (ActRatio) for comparison. The compression ratio $\gamma$ is set to 0.51.00.5-1.0, respectively.
Evaluation Metrics
We evaluate the performance of our method using three primary metrics: Accuracy (%), which represents the percentage of problems for which the final answer is correct. Tokens, the average number of tokens generated in the reasoning chain, measuring computational cost and efficiency during the reasoning process. ActRation, the actual compression ratio of ASCoT, reflecting the real efficiency of the token compression process.
Main Results
We evaluate the end-to-end performance of ASCoT against two prevalent CoT length control techniques. The first is Prompt, appends an explicit instruction (e.g., “Please reduce your CoT output by 50%”) to the model’s input, thus guiding the LLM to generate a shorter reasoning trace. The second baseline, termed Truncation, imposes a fixed upper bound on the number of output tokens during inference.
The results of LLaMA-3.1-8B-Instruct are summarized in Table 1. As shown, prompt‐based reduction consistently fails to meet the target ratio: even at a nominal 0.5 budget it yields an actual compression ratio of 0.94 on MATH-500, with only marginal savings in CoT tokens. Truncation strictly enforces the specified ratio but at the cost of dramatic performance degradation: at a budget of 0.5, the accuracy on GSM8K plummets by 79.2 points and by 21.2 points on MATH-500, rendering it unsuitable for tasks requiring robust multistep reasoning. In contrast, ASCoT closely matches the specified budget (e.g., ActRatio = 0.55 at Ratio = 0.5) while maintaining robust reasoning performance. In a full budget (Ratio = 1.0), ASCoT marginally outperforms the baseline, achieving 86.9% on GSM8K (+0.7) and 48.8% on MATH-500 (+0.2). In at a 0.5 budget, it maintains respectable accuracies of 79.5% (–6.7) on GSM8K and 40.1% (–8.5) on MATH-500, respectively. These results demonstrate that ASCoT strikes a favorable balance between compression and accuracy, validating its effectiveness.
We also provide the results of Qwen2.5‐Instruct series models as Table 2. Across the three model scales (3B, 7B, and 14B), ASCoT consistently maintains original performance while achieving substantial CoT compression. For the 3B model, ASCoT at 0.9 compression incurs only a 0.5 point drop in accuracy, while reducing token usage by 16% (314.87 $\rightarrow$ 262.31). As the compression ratio decreases to 0.5, the accuracy degrades from 83.5% to 74.9%, but the token count falls sharply to 170.87(reducing 45.7%). Similar trends hold for the 7B and 14B variants: at Ratio = 0.9, both scales suffer less than a 0.5 point accuracy drop (for the 14B model, there is even a slight improvement in performance), yet token counts shrink by about 15%. Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget.
Figure 4: Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
Scale Methods Ratio Accuracy Tokens ActRatio
3B Original 83.7(0.0↓) 314.87
ASCoT 1.0 83.5(0.2↓) 318.32 1.00
0.9 83.2(0.5↓) 262.31 0.82
0.8 82.0(1.7↓) 250.75 0.79
0.7 80.1(3.6↓) 233.15 0.73
0.6 77.6(6.1↓) 199.44 0.63
0.5 74.9(8.8↓) 170.87 0.54
7B Original 91.4(0.0↓) 297.83
ASCoT 1.0 91.8(0.4↑) 294.74 1.00
0.9 91.3(0.1↓) 251.71 0.85
0.8 90.4(1.0↓) 234.95 0.80
0.7 90.3(1.1↓) 216.73 0.74
0.6 88.2(3.2↓) 178.28 0.60
0.5 86.3(5.1↓) 150.98 0.51
14B Original 93.1(0.0↓) 313.11
ASCoT 1.0 93.3(0.2↑) 312.30 1.00
0.9 93.6(0.5↑) 269.76 0.86
0.8 93.5(0.4↑) 246.57 0.79
0.7 93.7(0.6↑) 221.17 0.71
0.6 92.8(0.3↓) 177.40 0.57
0.5 91.6(1.5↓) 158.21 0.51
Table 2: Experimental results on the Qwen2.5-Instruct series. We also provide Accuracy, Tokens and ActRatio.
Error Propagation via Controlled Injection
To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conduct a systematic error-injection study to test the “cascading failure” hypothesis. Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark. For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of correct CoT sequences. Then, we evaluate how these perturbations affect the overall reasoning process.
The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation. By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position of such errors affects the final reasoning outcome.
Numeric errors were introduced to simulate arithmetic miscalculations by slightly altering a single number in the reasoning process (e.g., adding or subtracting a fixed constant value). Symbolic errors were introduced to simulate logical inconsistencies by replacing a primary arithmetic operator with its inverse (e.g. + to -, $\times$ to $\div$). We preemptively set strict constraints to avoid division by zero.
Table 3 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B. When injecting a numerical error, the accuracy drops are relatively uniform across positions: such as a miscalculation at step 4 a drop of 61.70%. In contrast, symbolic errors induce more severe degradation, and this severity escalates with later injection points. Replacement of an operator in the final step (4/4) inflicts a reduction of 80.80%. These results confirm that CoT reasoning exhibits “late-stage fragility” particularly to logical perturbations. More details can be found in the appendix.
Analysis
Length Budget Analysis
To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget. As illustrated in Figure  4, we assess our approach under two distinct length-budget settings. In the first, the maximum generation length is scaled by the compression ratio (0.5–1.0). While in the second, we retain the original LLM budget (e.g., 1024 tokens in MATH) regardless of compression. The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%. Notably, at compression ratios of 0.8, and 0.9, we observe absolute gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT within the same budgetary constraints.
The Late-Stage Fragility of CoT
Our findings reveal a phenomenon that challenges the traditional cascading failure model. To illustrate this conceptually, Figure  5 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process.
An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps. However, an identical error introduced in the final or penultimate step was significantly more likely to go uncorrected, directly leading to a wrong answer.
Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it. Previous work suggests that early-stage errors in CoT reasoning tend to propagate forward, compromising subsequent steps and the final answer. However, our analysis reveals that the impact of reasoning errors is not uniform. Errors introduced near the conclusion of reasoning are notably much less likely to be corrected. This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
Figure 5: A illustration of “Late-Stage Fragility”. The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step. The tokens that occur related errors are highlighted in orange.
Case Study: ASCoT in Action
To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure  3. The problem asks for the total number of apples Bob picked over nine days.
As shown in the Initial CoT panel, a standard model correctly structures the reasoning but commits a calculation error in Step 2, incorrectly stating that 125 * 8 = 1,005. This single subtle yet critical flaw propagates directly to the final step, resulting in an incorrect answer of 1,020.
The Evaluation panel demonstrates the intervention of AVM module. The AVM performs a multi-dimensional assessment of the flawed step. It correctly identifies the lack of Logical Validity (0/1) due to the arithmetic error. Simultaneously, it recognizes that the step’s Factual Groundedness and Semantic Clarity & Process Utility are sound. This evaluation allows the AVM to pinpoint the exact failure without discarding the entire step’s valid structure.
Because the step’s risk score exceeds the threshold, the MSCE is activated. As shown in the Final result panel, the MSCE provides the corrected calculation (125 * 8 = 1,000) using a dual-path correction mechanism. This targeted repair allows the model to obtain the correct final answer of 1,015.
All/Err Type Ori Acc Fin Acc Drop Ratio
2/2 numeric 98.85 38.93 59.92
3/3 96.19 35.72 60.47
4/4 95.75 34.05 61.70
2/2 symbolic 98.85 29.18 69.67
3/3 96.19 20.42 75.77
4/4 95.75 14.95 80.80
Table 3: Error-injection results on GSM8K, using Qwen2.5-Instruct 7B. “All/Err” denotes the total number of reasoning steps and the step which a fault was injected. “Type” indicates the category of the injected error (numeric or symbolic). “Ori Acc” is the model’s accuracy before error-injection. “Fin Acc” is the accuracy after injection. “Drop Ratio” quantifies the relative decrease in accuracy.
Related Work
LLM Reasoning & Efficient Thinking
The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs (Marconato et al. 2023; Ding et al. 2024; Liu et al. 2024d). CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks (Team 2024).
These efficiency-focused methods can be broadly categorized: Prompt-based strategies (Xu et al. 2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget (Han et al. 2024). Inference-time dynamic reasoning (Ding et al. 2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation. CoT Distillation (Wadhwa, Amir, and Wallace 2024; Wang et al. 2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
CoT Compression and Pruning
The verbosity of CoT (Wingate, Shoeybi, and Sorensen 2022), while beneficial for reasoning, leads to high inference latency and computational costs. This has spurred research into CoT compression and pruning (Chuang et al. 2024; Xuefei et al. 2023), which aims to eliminate redundant information without degrading reasoning accuracy. This issue is often termed the “overthinking phenomenon,” (Kumar et al. 2025; Chen et al. 2024b) where models generate unnecessarily long chains that can even introduce errors.
To address this bottleneck, researchers have explored various techniques. Token-level pruning (Fu et al. 2024; Ma, Fang, and Wang 2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts. Other approaches operate at a more structural level. Prune-on-Logic (Liu et al. 2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score. Simpler methods, such as direct truncation of the reasoning chain (Wang et al. 2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
Self-correction Mechanism
The capacity for self-correction is a critical frontier (Pan et al. 2023) for enhancing the reliability of LLMs. The general paradigm (Kamoi et al. 2024; Yang et al. 2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
Approaches diverge based on the source of the feedback. Intrinsic self-correction (Li et al. 2024; Liu et al. 2024a) relies on the model’s own capabilities to critique and improve its output without external tools (Cao, Wang, and Xiong 2025; Tan et al. 2025). However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data (Liu et al. 2024b; Anantaprayoon, Kaneko, and Okazaki 2025). Extrinsic self-correction incorporates feedback from external sources. This involves using a verifier model to assess the correctness of each step (Kim et al. 2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
Conclusion
In this work, we challenge the conventional “cascading failure” hypothesis in CoT reasoning. Through systematic error-injection experiments, we identify and quantify “Late-Stage Fragility”: a counter-intuitive phenomenon where errors introduced near the end of a reasoning chain are significantly more detrimental than those at the beginning. To counteract this specific vulnerability, we introduce the Adaptive self-correction Chain-of-Thought (ASCoT) method, which is a targeted system that uses an adaptive verification manager to precisely identify and a multi-perspective engine to robustly correct these high-impact, late-stage errors. Our extensive experiments on the GSM8K and MATH benchmarks demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines including the computationally expensive self-consistency. These results highlight the effectiveness of ASCoT in addressing late-stage fragility, validating the importance of diagnosing specific failure modes to engineer more effective solutions. This research advocates for a shift from uniform verification to adaptive, position-aware strategies. We hope our investigations pave the way for future advancements in correction mechanisms and enhance the reliability, robustness, and efficiency of LLM reasoning.